OpenStudy (animelover98):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. This is my equation y=4(x-10)^2+27 Do I just solve for 4=4(x-10)^2+27 and 7=4(x-10)^2+27

OpenStudy (anonymous):

First plug in each of your x values (4 & 7) and solve for each of those. It will give you the corresponding y values. Then you are going to set up a second equation it will be the SECOND y values MINUS the FIRST y value over the SECOND x value MINUS the FIRST x value. It will give you a fraction and that is the rate of change.

OpenStudy (animelover98):

Oh ok, so it will be written as f(x)=4(4-10)^2+27 and f(x)=4(4-10)^2+27?

OpenStudy (animelover98):

I meant 7 on the second equation, sorry.

OpenStudy (anonymous):

Yeah, then you want to simplify.

OpenStudy (animelover98):

Thank you so much :) This was exactly the kind of help I needed.

OpenStudy (anonymous):

No problem :)